Country C produces 1,000 pounds of cotton at a cost of 50 cents per pound. They sell all of the cotton to Country S for 75 cents per pound. Country S makes 1,000 t-shirts with the cotton for a total cost of $1.50 per t-shirt. They sell all of the shirts to Country R for $2.00 each. Country R sells 950 of the t-shirts to domestic consumers for $10 each and the total cost of producing each shirt is $8 each. There are no other firms in this simple world. All income in the economy is either profit income or wage income 4. a. Find the value added by each country b. What is the value of world GDP? c. Percentage wise, how much income is earned by workers as opposed to the owners of capital in each country? prot 25o (332 00 1, 100

I understand most of this, but having trouble with C. Correct answers are shown, all I need to know is how to work through the problem. Please show work
Answer #1

First two are correct.

Country C produces 1000 pounds of Cotton at a cost of 50 cents per pound and sells at 75 cents per pound.

Here, 50 cent is the wage and 25 cent is the profit.

Wages of C = 0.5*1000 = $ 500

Next it is asked about owners capital not profit.

Country C owners capital = 50 cents

Wages of C = 100% of owners capital

Country S buys cotton at 75 cents and production cost = $ 1.50 pet t-shirt

Wage of S = 1.5 - 0.75 = $ 0.75 per t-shirt

Number of t-shirts= 1000

Total wages =$ 750

Owners capital = $ 1.50*1000 = $ 1500

Wages of S = 50% of owners capital

Country R: cost of producing shirt = $ 8

R purchased t-shirt at a price of $2

Wages of R = $ 6 each t shirt

Wages of R = (6/8)*100 = 75%

Question is about owners capital not profit.

Please contact if having any query thank you.